THE ROTATIONAL SPECTRUM OF CN^{−}

\underline{C. A. GOTTLIEB}, S. BRÜNKEN, M. C. McCARTHY, and P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138.

CN^{−} is a unique molecular anion because neutral CN has one of the highest known electron affinities, is observed under a wide range of terrestrial conditions, and because this closed-shell anion is structurally and spectroscopically quite simple, making it amenable to both laboratory detection and high-level quantum chemical calculations. Although CN^{−} has been studied by photoelectron spectroscopy,^{a} no high-resolution spectra have been reported. Here we report the laboratory rotational spectrum of CN^{−} observed in a low pressure glow discharge through C_{2}N_{2} and N_{2}. Determination of the rotational and centrifugal distortion constants and the nitrogen quadrupole hyperfine coupling constant allow line frequencies to be calculated well into the far IR. CN^{−} is an excellent candidate for astronomical detection, because the CN radical is observed in many galactic molecular sources, and calculations^{b} indicate CN^{−} should be detectable in IRC+10216 — the carbon star where C_{6}H^{−} has recently been observed. The fairly high concentration of CN^{−} in our discharge implies that other molecular anions containing the nitrile group may be within reach.

aS. E. Bradforth, E. H. Kim, D. W. Arnold, and D. M. Neumark, J. Chem. Phys. 98, 800 (1993).

bS. Petrie, Mon. Not. R. Astron. Soc. 281, 137 (1996).